OpenStudy (anonymous):

Solve the proportion below for X. X/17 = 3/5

OpenStudy (anonymous):

I will fan and give a medal for the best/right answer!!

OpenStudy (anonymous):

@jim_thompson5910 @phi

jimthompson5910 (jim_thompson5910):

Hint: cross multiply to go from X/17 = 3/5 to 5*x = 17*3 5x = 51 I'll let you finish

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

x = 10.2??

jimthompson5910 (jim_thompson5910):

that's your answer in decimal form the answer in fraction form is just 51/5

jimthompson5910 (jim_thompson5910):

so either x = 51/2 or x = 10.2 is your answer (depending on which form your teacher wants)

OpenStudy (anonymous):

I gave both just in case! Thank you!!

jimthompson5910 (jim_thompson5910):

you're welcome